# Enhancing the Efficiency of Time-Dependent Density Functional Theory Calculations of Dynamic Response Properties

### v1

@@ Abstract
X-ray Thomson scattering (XRTS) constitutes an essential technique for diagnosing material properties under extreme conditions, such as high pressures and intense laser heating. Time-dependent density functional theory (TDDFT) is one of the most accurate available ab initio methods for modeling XRTS spectra, as well as a host of other dynamic material properties. However, strong thermal excitations, along with the need to account for variations in temperature and density as well as the finite size of the detector significantly increase the computational cost of TDDFT simulations compared to ambient conditions. In this work, we present a broadly applicable method for optimizing and enhancing the efficiency of TDDFT calculations. Our approach is based on a one-to-one mapping between the dynamic structure factor and the imaginary time density–density correlation function, which naturally emerges in Feynman’s path integral formulation of quantum many-body theory. Specifically, we combine rigorous convergence tests in the imaginary time domain with a constraints-based noise attenuation technique to improve the efficiency of TDDFT modeling without the introduction of any significant bias. As a result, we can report a speed-up by up to an order of magnitude, thus potentially saving millions of CPU hours for modeling a single XRTS measurement of matter under extreme conditions.
@@ 1 Introduction
The advent of high brilliance 1, high resolution 2, 3, 4, 5, and high repetition rate 6 X-ray free-electron lasers (XFELs) such as the European XFEL (Germany) 7, 8, 9, LCLS (USA) 10, 1, and SACLA (Japan) 11 has significantly improved our ability to study the electronic structure properties of materials under both ambient and extreme conditions 12. Notable recent examples of successful XFEL applications in materials science under high pressures and/or high temperatures include the first experimental observations of liquid carbon 13, superheated gold crystal lattices beyond the predicted entropy catastrophe threshold 14, and phonon hardening in laser-excited gold 15. Moreover, modern backlighter set-ups 16 at high-power laser facilities such as the National Ignition Facility (NIF) 17 and the OMEGA laser facility 18 (both USA) have, despite their comparably lower spatial and temporal resolution, also become of considerable value for the diagnostics of extreme states of matter, with the recent study of the onset of pressure-driven K-shell delocalization in warm dense beryllium at the NIF 19 being a case in point.
@@
From the range of available X-ray diagnostics options, X-ray Thomson scattering (XRTS) has emerged as a primary technique for probing the state of matter under extreme conditions 20, 19, 21, 22. The resulting need to analyse high-resolution XRTS measurements has highlighted the importance of a true ab initio modeling of the probed electron dynamic structure factor (DSF) $S(\mathbf{q},\omega)$ 23. In addition, high-resolution XRTS data is not only used for diagnostics, but also serves as an important benchmark for theoretical models and to gain insights into the dielectric properties of materials 24, 25, 26, 27, 19, 4, 5. For example, identifying the applicability range of where common models are accurate is required for the creation of reliable, efficient, and flexible parameterized models of the dynamic collision frequency, electrical conductivity, and other material response properties 28, 29, 30. Such models, e.g., constitute critical input for state-of-the-art radiation hydrodynamics simulations 31, which are particularly important for the development of inertial fusion energy (IFE) applications; a field which is currently experiencing significant advancements 32, 33, 34 spearheaded by the recent demonstration of target gain greater than unity in an inertial fusion experiment at NIF35.
@@
The primary first-principles modeling method to compute the electronic DSF is given by time-dependent density functional theory (TDDFT). Both commonly used flavours—linear-response TDDFT 36, 37, 38, 39 and real-time TDDFT 40, 41, 42—utilize a Lorentzian or Gaussian broadening with a finite width $\eta$ to suppress numerical noise in the calculation of electron dynamic response functions and DSFs. This broadening often obscures important features of the DSF, making them difficult to discern. Theoretically, the correct limit for any response property would be achieved by setting the broadening parameter to $\eta\to 0$. However, in practice, $\eta$ must be set to a finite value to extract sufficiently smooth curves. For fixed simulation parameters, a decrease in $\eta$ results in the amplification of irregularities, which is characterized by evident ringing in the curves. Consequently, to achieve a smooth curve with lower $\eta$ values, a finer k-point mesh for sampling the first Brillouin zone is required. First and foremost, this presents a significant bottleneck for TDDFT as the computational cost increases cubically or even faster with the density of $k$-point mesh, coupled with the already inherently high computational demands of TDDFT. Secondly, the high computational cost of the TDDFT significantly obstructs the convergence tests for the DSF and its related dynamic response properties with respect to the parameter $\eta$.
@@
While ideal crystal structures exhibit lattice symmetries that help to mitigate these issues 43, these problems become much more severe in systems with strong thermal excitations due to a significant increase in the number of partially occupied bands as the temperature rises 44, 45. Moreover, for systems with disordered atomic structures such as disordered alloys, amorphous materials, and shock-compressed materials, techniques that employ crystal symmetries are not applicable . Additionally, for such materials, the computational cost increases dramatically due to the necessity of averaging the TDDFT results over various ionic configurations 46. Consequently, the optimal value for the broadening parameter $\eta$ is usually selected by visually assessing the smoothness of the TDDFT result for any given dynamic response property of interest 47, 48, 49; this is clearly a highly unsatisfactory approach for what is supposed to be a true first-principles calculation.
@@
In this work, we develop a novel metric to test the convergence with respect to $\eta$ that is formulated in the imaginary time domain, and combine it with a constraints-based noise attenuation technique, eliminating the need to continuously increase the density of $k$-points as $\eta$ decreases without the introduction of any significant bias. The cornerstone of our approach is the one-to-one mapping between the DSF $S(\mathbf{q},\omega)$ and the imaginary-time correlation density–density function (ITCF) $F(\mathbf{q},\tau)$—a key quantity both for the interpretation of XRTS experiments 21, 50, 23, 51, 52, 53, 54, 22, 27 and, e.g., for first-principles quantum Monte Carlo simulations 55, 56, 57, 58, 59, 60.
@@
The presented techniques boost the efficiency of ab initio TDDFT calculations of dynamic response properties for both ambient and extreme conditions, potentially allowing one to save on the order of millions of CPU hours for the modeling of XRTS spectra at extreme temperatures. In addition, the combination of the $\eta$-convergence test with the constraints-based noise attenuation enables more efficient TDDFT modeling of a plethora of other dielectric properties, such as the dynamic density response function, dielectric function, conductivity, absorption, etc. Therefore, we expect our work to become a new standard tool for the estimation of material properties and for the theoretical support of experimental measurements over a range of conditions and research fields. We demonstrate our new approach by applying it to the DSF of solid density hydrogen at parameters relevant for laser-heated hydrogen jet experiments 61, 62 and IFE 63, 64. Additionally, we consider aluminum (Al) both under ambient conditions and with isochoric heating 15, 2, 65, 66, 67.
@@
The paper is organized as follows: In Sec. 2, we present a method for the $\eta$-convergence test of TDDFT results in the imaginary time domain. In Sec. 3, we introduce a constraints-based noise attenuation method for TDDFT modeling of the DSF and other related dynamic response properties. In Sec. 4, we demonstrate the utility of the developed $\eta$-convergence test in the imaginary time domain combined with the constraints-based noise attenuation technique by analyzing the DSF of solid density hydrogen and isochorically heated aluminum. We conclude the paper by emphasizing our main findings and discussing potential applications. Supporting technical information about TDDFT calculations and data post-processing can be found in Appendix A.
@@ 2 Method for the $\eta$-convergence test in the imaginary time domain
Our proposed approach is based on the one-to-one mapping between the DSF $S(\mathbf{q},\omega)$ and the ITCF $F(\mathbf{q},\tau)$, indicating that both contain precisely the same information about dynamic material properties, only in different representations 68, 58, 69, 60. These two representations are connected through a two-sided Laplace transform:
@@
where $i\hbar\tau\in-i\hbar[0,\beta]$ is the imaginary time argument in the range defined by the inverse temperature $\beta=1/(k_{B}T)$, and the DSF at negative frequencies is computed using the detailed balance $S(\mathbf{q},-\omega)=S(\mathbf{q},\omega)\exp(-\beta\omega)$; the latter is equivalent to the symmetry relation of the ITCF $F(\mathbf{q},\tau)=F(\mathbf{q},\beta-\tau)$, which holds universally in thermal equilibrium 21, 50. The ITCF naturally emerges in Feynman’s path integral formulation of quantum many-body theory. Recently, it has gained increasing attention after Dornheim et al. 21, 50 demonstrated that the two-sided Laplace transform allows for the direct determination of the electron temperature, and a variety of other properties 23, 53, 22, 54, 27, from XRTS measurements without the need for any modeling.
For the current analysis, an important aspect of the Laplace transform is its ability to filter out high-frequency noise 50, making it a robust tool for performing the $\eta$-convergence test. To demonstrate this point, we show the results for the DSF and corresponding ITCF computed for solid density hydrogen with temperature $T=4.8~{\rm eV}$ and density $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$ and the Lorentzian smearing parameters $\eta=0.05~{\rm eV}$, $\eta=0.1~{\rm eV}$, and $\eta=0.5~{\rm eV}$ in Fig. 1. At this momentum transfer value, the DSF is primarily determined by collective plasmon scattering. Panel (a) clearly illustrates the ambiguity and difficulty of assessing the $\eta$-convergence of the DSF at a given set of simulation parameters (cutoff energy, number of bands, $k$-point grid, etc.). Increasing $\eta$ from $0.05~{\rm eV}$ to $0.1~{\rm eV}$ significantly reduces the magnitude of fluctuations in the DSF. The results computed at $\eta=0.05~{\rm eV}$ effectively encompass those generated at $\eta=0.1~{\rm eV}$. When $\eta$ is further increased to $0.5~{\rm eV}$, we obtain a much smoother curve. This curve remains within the range covered by the data at $\eta=0.05~{\rm eV}$, but it shows reduced values around the plasmon region and in the low-frequency range ($\omega\lesssim 15~{\rm eV}$) to the degree that it barely overlaps with the results at $\eta=0.1~{\rm eV}$. In contrast, as we show in Fig. 1(b), the ITCFs computed using $\eta=0.05~{\rm eV}$ and $\eta=0.1~{\rm eV}$ are in excellent agreement with each other, while the ITCF calculated using $\eta=0.5~{\rm eV}$ significantly deviates from both of them. This exemplifies the suitability of the ITCF for the $\eta$-convergence test as the smoother DSF at $\eta=0.5~{\rm eV}$ comes at the cost of a significant unphysical bias that can be clearly detected in $F(\mathbf{q},\tau)$.
@@
This adjustment is necessary because TDDFT data for a fixed configuration of ions do not capture the quasi-elastic part of the DSF 70, 71 (e.g., caused by phonons or ionic plasma oscillations), which translates to a constant shift of the ITCF that does not depend on $\tau$ 72. Also, TDDFT does not allow for the adequate computation of the $\omega\to 0$ limit of the DSF due to finite size effects 73. A similar effect is well known, e.g., for the dynamic Onsager coefficients from the Kubo-Greenwood method 74. The impact of these finite-size effects at the $\omega\to 0$ limit of the DSF is also manifested as an uncertainty in a constant shift in $F(\mathbf{q},\tau)$ 73.
@@
We present a more detailed and systematic $\eta$-convergence test of the ITCF for solid density hydrogen at a temperature of $T=4.8~{\rm eV}$ and a momentum transfer of $q=0.946~\textup{~\AA }^{-1}$ in Fig. 2. In panel (a), we show the dependence of the ITCF minimum, $F(\mathbf{q},\tau=\beta/2)$, on $\eta$. As stated above, this minimum directly follows from the detailed balance relation in thermal equilibrium. Additionally, we display the $\eta$-dependence of the area under the ITCF, which corresponds to the static electronic density response scaled by a constant factor 58, $\chi(\mathbf{q})=-n_{e}\int_{0}^{\beta}F(\mathbf{q},\tau)\,{\mathrm{d}}\tau$. We observe that the minimum of the ITCF and the area under the ITCF both exhibit a weak dependence on $\eta$ for $\eta\leq 0.1~{\rm eV}$, while there is a strong dependence for $\eta>0.1~{\rm eV}$. Due to the previously mentioned finite-size effects in the DSF at small frequencies, there is a small reduction in both the ITCF minimum and the area under the ITCF as $\eta$ decreases below $0.05~{\rm eV}$. Therefore, we consider $\widetilde{F}(\mathbf{q},\tau=\beta/2)$ and $\int_{0}^{\beta}\widetilde{F}(\mathbf{q},\tau)\,{\mathrm{d}}\tau$ in Fig. 2(b), as they are not afflicted by these artifacts and clearly exhibit convergence with respect to $\eta$ at $\eta\leq 0.1~{\rm eV}$. From this, we draw the important intermediate conclusion that TDDFT calculations for the ITCF can achieve convergence with respect to the parameter $\eta$. Therefore, the largest $\eta$ within the convergence range provides the DSF that exhibits the smallest magnitude of numerical noise for a given set of simulation parameters without showing any detectable bias in the Laplace transform, and the corresponding DSF data is referred to as the reference DSF.
@@
We emphasize that the generation of the DSFs with various $\eta$ takes place during the post-processing phase of the TDDFT simulation75, 40, 41 and requires negligible computational time. Therefore, the search for the reference DSF using the $\eta$-convergence test in the imaginary time domain adds negligible computational overhead. To further reduce the noise level in the DSF, there now are two options: (i) one might increase the number of $k$-points and/or the system size, which are both computationally demanding and, indeed, can even be prohibitive; here, we advocate for a second alternative strategy (ii) in using a constraints-based noise-filter that removes the artificial ringing from the DSF without the need for additional TDDFT efforts, and without the introduction of any significant bias as it occurs in the $\eta$-smearing for larger values of $\eta$.
3 Constraints-based noise attenuation

### v3

@@ Abstract
X-ray Thomson scattering (XRTS) constitutes an essential technique for diagnosing material properties under extreme conditions, such as high pressures and intense laser heating. Time-dependent density functional theory (TDDFT) is one of the most accurate available ab initio methods for modeling XRTS spectra, as well as a host of other dynamic material properties. However, strong thermal excitations, along with the need to account for variations in temperature and density as well as the finite size of the detector significantly increase the computational cost of TDDFT simulations compared to ambient conditions. In this work, we present a broadly applicable method for optimizing and enhancing the efficiency of TDDFT calculations. Our approach is based on a one-to-one mapping between the dynamic structure factor and the imaginary time density–density correlation function, which naturally emerges in Feynman’s path integral formulation of quantum many-body theory. Specifically, we combine rigorous convergence tests in the imaginary time domain with a constraints-based attenuation of narrow-band fluctuations to improve the efficiency of TDDFT modeling without the introduction of any significant bias. As a result, we can report a speed-up by up to an order of magnitude, thus substantially reducing the burden of computational cost required for XRTS analysis.
@@ 1 Introduction
The advent of high brilliance Fletcher2015, high resolution Descamps2020, Wollenweber_RSI, Gawne_PRB_2024, Gawne_2025, and high repetition rate Tschentscher_2017 X-ray free-electron lasers (XFELs) such as the European XFEL (Germany) preston2020design, Zastrau_2021, Descamps_RSI_2025, LCLS (USA) Ofori_Okai_2018, Fletcher2015, and SACLA (Japan) hara2013two has significantly improved our ability to study the electronic structure properties of materials under both ambient and extreme conditions Pascarelli2023. Notable recent examples of successful XFEL applications in materials science under high pressures and/or high temperatures include the first experimental observations of liquid carbon Kraus_Nature_2025, superheated gold crystal lattices beyond the predicted entropy catastrophe threshold White_Nature_2025, and phonon hardening in laser-excited gold Descamps_sciadv. Moreover, modern backlighter set-ups MacDonald_POP_2022 at high-power laser facilities such as the National Ignition Facility (NIF) Moses_NIF and the OMEGA laser facility Glenzer_PRL_2007 (both USA) have, despite their comparably lower spatial and temporal resolution, also become of considerable value for the diagnostics of extreme states of matter, with the recent study of the onset of pressure-driven K-shell delocalization in warm dense beryllium at the NIF Tilo_Nature_2023 being a case in point.
@@
From the range of available X-ray diagnostics options, X-ray Thomson scattering (XRTS) has emerged as a primary technique for probing the state of matter under extreme conditions siegfried_review, Tilo_Nature_2023, Dornheim_T_20222, Bellenbaum_APL_2025. The resulting need to analyse high-resolution XRTS measurements has highlighted the importance of a true ab initio modeling of the probed electron dynamic structure factor (DSF) $S(\mathbf{q},\omega)$ dornheim2024_Rayleigh. In addition, high-resolution XRTS data is not only used for diagnostics, but also serves as an important benchmark for theoretical models and to gain insights into the dielectric properties of materials Sperling_PRL_2015, Sperling_2017, Witte_PRL, Dornheim_NatureCom_2025, Tilo_Nature_2023, Gawne_PRB_2024, Gawne_2025. For example, identifying the applicability range of where common models are accurate is required for the creation of reliable, efficient, and flexible parameterized models of the dynamic collision frequency, electrical conductivity, and other material response properties Hentschel_POP_2025, Hentschel_POP_2023, Moldabekov_PPNP_2025. Such models, e.g., constitute critical input for state-of-the-art radiation hydrodynamics simulations Clark_POP_2015, which are particularly important for the development of inertial fusion energy (IFE) applications; a field which is currently experiencing significant advancements Batani_roadmap, Spears_Science_2025, Betti2023 spearheaded by the recent demonstration of target gain greater than unity in an inertial fusion experiment at NIFPhysRevLett.132.065102.
@@
The primary first-principles modeling method to compute the electronic DSF is given by time-dependent density functional theory (TDDFT). Both commonly used flavours—linear-response TDDFT book_Ullrich, Byun_2020, Moldabekov_PRR_2023, Moldabekov_mre_2025 and real-time TDDFT Baczewski_PRL_2016, White_Electronic_Structure_2025, Dornheim_review—utilize a Lorentzian or Gaussian broadening with a finite width $\eta$ to suppress numerical fluctuations in the results for electron dynamic response functions and DSFs. This broadening often obscures important features of the DSF, making them difficult to discern. Theoretically, the correct limit for any response property would be achieved by setting the broadening parameter to $\eta\to 0$. However, in practice, $\eta$ must be set to a finite value to extract sufficiently smooth curves for DSFs. For fixed simulation parameters, a decrease in $\eta$ results in the amplification of irregularities, which is characterized by evident ringing in the curves. Consequently, to achieve a smooth curve with lower $\eta$ values, a finer k-point mesh for sampling the first Brillouin zone is required. First and foremost, this presents a significant bottleneck for TDDFT as the computational cost increases cubically or even faster with the density of $k$-point mesh, coupled with the already inherently high computational demands of TDDFT. Secondly, the high computational cost of the TDDFT significantly obstructs the convergence tests for the DSF and its related dynamic response properties with respect to the parameter $\eta$.
@@
While ideal crystal structures exhibit lattice symmetries that help to mitigate these issues MacDonald_1979, these problems become much more severe in systems with strong thermal excitations due to a significant increase in the number of partially occupied bands as the temperature rises Blanchet_POP_2020, Fiedler_PRR_2022. Moreover, for systems with disordered atomic structures such as disordered alloys, amorphous materials, and shock-compressed materials, techniques that employ crystal symmetries are not applicable. Additionally, for such materials, the computational cost increases dramatically due to the necessity of averaging the TDDFT results over various ionic configurations JCP_averaging. Consequently, the optimal value of the broadening parameter $\eta$ in such computationally demanding cases is usually selected by visually assessing the smoothness of the TDDFT results for the dynamic response property of interest abinit_tddft, Meghan_AIP_2023, Desjarlais_PRE_2002; this is clearly a highly unsatisfactory approach for what is supposed to be a true first-principles calculation.
@@
In this work, we develop a novel metric to test the convergence with respect to $\eta$ that is formulated in the imaginary time domain, and combine it with a constraints-based attenuation of narrow-band fluctuations, eliminating the need to continuously increase the density of $k$-points as $\eta$ decreases without the introduction of any significant bias. The cornerstone of our approach is the one-to-one mapping between the DSF $S(\mathbf{q},\omega)$ and the imaginary-time correlation density–density function (ITCF) $F(\mathbf{q},\tau)$—a key quantity both for the interpretation of XRTS experiments Dornheim_T_20222, Dornheim_T2_2022, dornheim2024_Rayleigh, Schoerner_PRE_2023, shi2025firstprinciplesanalysiswarmdense, schwalbe2025staticlineardensityresponse, Vorberger_PRX_2023, Bellenbaum_APL_2025, Dornheim_NatureCom_2025 and, e.g., for first-principles quantum Monte Carlo simulations boninsegni1996, Dornheim_JCP_ITCF_2021, Saccani_Supersolid_PRL_2012, Dornheim_mre_2023, Dornheim_MRE_2024, Chuna_JCP_2025.
@@
The presented techniques boost the efficiency of ab initio TDDFT calculations of dynamic response properties for both ambient and extreme conditions, thereby substantially reducing the computational effort required for modeling XRTS spectra. In addition, the combination of the $\eta$-convergence test with the constraints-based attenuation of narrow-band fluctuations enables more efficient TDDFT modeling of a plethora of other dielectric properties, such as the dynamic density response function, dielectric function, conductivity, absorption, etc. Therefore, we expect our work to become a new standard tool for the estimation of material properties and for the theoretical support of experimental measurements over a range of conditions and research fields. We demonstrate our new approach by applying it to the DSF of solid density hydrogen (with $\rho=0.08~{\rm g/cc}$) at parameters relevant for experiments with laser heating of cryogenic hydrogen targets Fletcher_Frontiers_2022, Zastrau and IFE Hu_PRL_2010, Bonitz_pop_2024. Additionally, we consider aluminum (Al) both under ambient conditions and with isochoric heating Descamps_sciadv, Descamps2020, Moldabekov_prr_2024, Moldabekov_Omega_2024, Sperling_2015.
@@
The paper is organized as follows: In Sec. 2, we present a method for the $\eta$-convergence test of TDDFT results in the imaginary time domain. In Sec. 3, we introduce a method of constraints-based attenuation of narrow-band fluctuations for TDDFT modeling of the DSF and other related dynamic response properties. In Sec. 4, we demonstrate the utility of the developed $\eta$-convergence test in the imaginary time domain combined with the constraints-based attenuation of narrow-band fluctuations by analyzing the DSF of solid density hydrogen and isochorically heated aluminum. We conclude the paper by emphasizing our main findings and discussing potential applications. Supporting technical information about TDDFT calculations and data post-processing can be found in Appendix A.
@@ 2 Method for the $\eta$-convergence test in the imaginary time domain
Our proposed approach is based on the one-to-one mapping between the DSF $S(\mathbf{q},\omega)$ and the ITCF $F(\mathbf{q},\tau)$, indicating that both contain precisely the same information about dynamic material properties, only in different representations Dornheim_prb_2023, Dornheim_mre_2023, Dornheim_PTR_2022, Chuna_JCP_2025. These two representations are connected through a two-sided Laplace transform:
@@
where $i\hbar\tau\in-i\hbar[0,\beta]$ is the imaginary time argument in the range defined by the inverse temperature $\beta=1/(k_{B}T)$, and the DSF at negative frequencies is computed using the detailed balance $S(\mathbf{q},-\omega)=S(\mathbf{q},\omega)\exp(-\beta\omega)$; the latter is equivalent to the symmetry relation of the ITCF $F(\mathbf{q},\tau)=F(\mathbf{q},\beta-\tau)$, which holds universally in thermal equilibrium Dornheim_T_20222, Dornheim_T2_2022. The ITCF naturally emerges in Feynman’s path integral formulation of quantum many-body theory. Recently, it has gained increasing attention after Dornheim et al. Dornheim_T_20222, Dornheim_T2_2022 demonstrated that the two-sided Laplace transform allows for the direct determination of the electron temperature, and a variety of other properties dornheim2024_Rayleigh, schwalbe2025staticlineardensityresponse, Bellenbaum_APL_2025, Vorberger_PRX_2023, Dornheim_NatureCom_2025, from XRTS measurements without the need for any modeling.
For the current analysis, an important aspect of the Laplace transform is its ability to filter out high-frequency narrow-band fluctuations Dornheim_T2_2022, making it a robust tool for performing the $\eta$-convergence test. To demonstrate this point, we show the results for the DSF and corresponding ITCF computed for solid density hydrogen with temperature $T=4.8~{\rm eV}$ and density $\rho=0.08~{\rm g/cc}$ at $q=0.946~\textup{~\AA }^{-1}$ and the Lorentzian smearing parameters $\eta=0.05~{\rm eV}$, $\eta=0.1~{\rm eV}$, and $\eta=0.5~{\rm eV}$ in Fig. 1. At this momentum transfer value, the DSF is primarily determined by collective plasmon scattering. Panel (a) clearly illustrates the ambiguity and difficulty of assessing the $\eta$-convergence of the DSF at a given set of simulation parameters (cutoff energy, number of bands, $k$-point grid, etc.). Increasing $\eta$ from $0.05~{\rm eV}$ to $0.1~{\rm eV}$ significantly reduces the magnitude of fluctuations in the DSF. The results computed at $\eta=0.05~{\rm eV}$ effectively encompass those generated at $\eta=0.1~{\rm eV}$. When $\eta$ is further increased to $0.5~{\rm eV}$, we obtain a much smoother curve. This curve remains within the range covered by the data at $\eta=0.05~{\rm eV}$, but it shows reduced values around the plasmon region and in the low-frequency range ($\omega\lesssim 15~{\rm eV}$) to the degree that it barely overlaps with the results at $\eta=0.1~{\rm eV}$. In contrast, as we show in Fig. 1(b), the ITCFs computed using $\eta=0.05~{\rm eV}$ and $\eta=0.1~{\rm eV}$ are in excellent agreement with each other, while the ITCF calculated using $\eta=0.5~{\rm eV}$ significantly deviates from both of them. This exemplifies the suitability of the ITCF for the $\eta$-convergence test as the smoother DSF at $\eta=0.5~{\rm eV}$ comes at the cost of a significant unphysical bias that can be clearly detected in $F(\mathbf{q},\tau)$.
@@
This adjustment is necessary because TDDFT data for a fixed configuration of ions do not capture the quasi-elastic part of the DSF Mo_prb_2020, Gregori_PRE_2003 (e.g., caused by phonons or ionic plasma oscillations), which translates to a constant shift of the ITCF that does not depend on $\tau$ Bellenbaum_prr_2025. Also, TDDFT does not allow for the adequate computation of the $\omega\to 0$ limit of the DSF due to finite size effects moldabekov2025abinitiodensityfunctional. A similar effect is well known, e.g., for the dynamic Onsager coefficients from the Kubo-Greenwood method Melton_POP_2024. The impact of these finite-size effects at the $\omega\to 0$ limit of the DSF is also manifested as an uncertainty in a constant shift in $F(\mathbf{q},\tau)$ moldabekov2025abinitiodensityfunctional.
@@
We present a more detailed and systematic $\eta$-convergence test of the ITCF for solid density hydrogen at a temperature of $T=4.8~{\rm eV}$ and a momentum transfer of $q=0.946~\textup{~\AA }^{-1}$ in Fig. 2. In panel (a), we show the dependence of the ITCF minimum, $F(\mathbf{q},\tau=\beta/2)$, on $\eta$. As stated above, this minimum directly follows from the detailed balance relation in thermal equilibrium. Additionally, we display the $\eta$-dependence of the area under the ITCF, which corresponds to the static electronic density response scaled by a constant factor Dornheim_mre_2023, $\chi(\mathbf{q})=-n_{e}\int_{0}^{\beta}F(\mathbf{q},\tau)\,{\mathrm{d}}\tau$. We observe that the minimum of the ITCF and the area under the ITCF both exhibit a weak dependence on $\eta$ for $\eta\leq 0.1~{\rm eV}$, while there is a strong dependence for $\eta>0.1~{\rm eV}$. Due to the previously mentioned finite-size effects in the DSF at small frequencies, there is a small reduction in both the ITCF minimum and the area under the ITCF as $\eta$ decreases below $0.05~{\rm eV}$. Therefore, we consider $\widetilde{F}(\mathbf{q},\tau=\beta/2)$ and $\int_{0}^{\beta}\widetilde{F}(\mathbf{q},\tau)\,{\mathrm{d}}\tau$ in Fig. 2(b), as they are not afflicted by these artifacts and clearly exhibit convergence with respect to $\eta$ at $\eta\leq 0.1~{\rm eV}$. From this, we draw the important intermediate conclusion that TDDFT calculations for the ITCF can achieve convergence with respect to the parameter $\eta$. Therefore, the largest $\eta$ within the convergence range provides the DSF that exhibits the smallest magnitude of numerical fluctuations for a given set of simulation parameters without showing any detectable bias in the Laplace transform, and the corresponding DSF data is referred to as the reference DSF.
@@
We emphasize that the generation of the DSFs with various $\eta$ takes place during the post-processing phase of the TDDFT simulationTIMROV2015460, Baczewski_PRL_2016, White_Electronic_Structure_2025 and requires negligible computational time. Therefore, the search for the reference DSF using the $\eta$-convergence test in the imaginary time domain adds negligible computational overhead. To further reduce the level of non-physical fluctuations in the DSF, there now are two options: (i) one might increase the number of $k$-points and/or the system size, which are both computationally demanding and, indeed, can even be prohibitive; here, we advocate for a second alternative strategy (ii) in using a constraints-based filter that removes the artificial ringing from the DSF without the need for additional TDDFT efforts, and without the introduction of any significant bias as it occurs in the $\eta$-smearing for larger values of $\eta$.
3 Constraints-based attenuation of narrow-band fluctuations